OWG Meeting 23-Apr-2014
Location and Time
Room: CC F326-327
(if problems, call phone in conference room: 757-269-6460)
- To join via Polycom room system go to the IP Address: 18.104.22.168 (bjn.vc) and enter the meeting ID: 120390084.
- To join via a Web Browser, go to the page https://bluejeans.com/120390084.
- To join via phone, use one of the following numbers and the Conference ID: 120390084
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- More information on connecting to bluejeans is available.
- Collaboration Meeting May 12-14
- Major systems reports
- DAQ system Configuration
- Mantis Task Tracker
David L., Bryan M., Sergey F., Simon T., Alex S., HOvanes E., Dmitry R., Beni Z., Vanik, Serguei P., Curtis M., Eugene C., Mark I., Michael S., Dave A., Vardan G., Nerses G.
- Fiber (for networking) is installed
- CC will connect to switches.
- Can do this without disrupting network service (in principle)
- Preliminarily planned for May 5th.
- Wiki update is scheduled for this evening
- New detector-specific accounts have been created on gluon cluster
- Working on BCAL Cosmic Trigger
- 3 crates (1 BCAL and 2 trigger crates)
- Pulser feeding into fADC modules in BCAL crate
- Have done one test with entire chain (trigger originating from BCAL and going through TS back to BCAL to trigger event)
- Revised the front-end configuration to use 3 files
- One is the global level settings
- Second is crate-level configurations
- Third is optional user supplied configurations that can be used to mask channels or disable entire boards
- Will implement read back of register values to confirm they were correctly set
- 5 flash modules changed in FCAL
- FCAL has noticed 3 failure modes in individual fADC250 channels (not always same channels)
- pedestal shifted by some amount (bad DAC setting?)
- time shifted by some fixed amount
- channel value locks at single value.
- Not yet verifying DAC values set are actually set. (will do soon)
- Bryan Moffit noted that Hai has narrowed down the location of one of the problems to a small portion of the FPGA code. They may even have a fix.
- Bryan would like to look at the ROL to confirm the initialization sequencing so that can be eliminated as a possible cause.
- One of the TOF crates is missing the TI module (Alex thinks it's in his office )
- TOF DAQ:
- Someone needs to take responsibility for the TOF dedicated DAQ (Sergey F. was nominated twice for this role)
- Some discussion about using TOF as cosmic trigger for other detector systems
- The trigger algorithm for the TOF is not ready and not being worked at the moment. It is lower on priority list
- May be possible to use BCAL simple sum algorithm for trigger
- Work continues, but nothing significant to report
- Nerses is working on BCAL GUI problems
- Vanik is here and working on motor applications
- Dave Butler continues work on gas system
- Mike S. working on EPICS guis that will be used by shift personnel
- Orlando is working on scripts for running pulser sequence for BCAL
- Accelerator group is working on developing some guis for us to control magnets and other beam line devices
- CDC voltage controls are now operational
DAQ system Configuration
An open discussion was had on some of the technical details on how to organize the configuration information for the DAQ system and related components. The Trigger Group has developed a nearly complete system. It could incorporate other items (e.g. monitoring) but that would couple the systems. Several ideas were discussed though unanimous consent was not reached on any particular plan. One significant agreement was that we can use a file-based config system for all detectors providing it is coupled to a database system that provides a full history mechanism.